OpenStudy (anonymous):

What is the solution set of the following equation? -8x+x+15=-7x+12 a) Ø b) {1/7} c) {all reals}

OpenStudy (anonymous):

option c) {all reals}

OpenStudy (anonymous):

-8x + x + 15 = -7x + 12 -7x + 15 = -7x + 12 -7x + 7x = 12 - 15 0 = -3 (incorrect) This equation has no solution

OpenStudy (anonymous):

Thank you, could you explain how you g the answer please? I have no idea how to do these problems.

OpenStudy (anonymous):

ya @kelliegirl33 you are right...I am sorry

OpenStudy (anonymous):

-8x + x + 15 = -7x + 12 -- start by combining like terms -7x + 15 = -7x + 12 -- add 7x to both sides -7x + 7x + 15 = -7x + 7x + 12 15 = 12 (not true) if you get an answer that is not true, then there is no solution

OpenStudy (anonymous):

Thank you!